In this addition probleM, each letter represents the same digit in both problem. Replace each letter with a different digit, 1 through 9, so that both additon problems a true.
ABC + DEF = GHI

ADG + BEH = CFI

Answer by Edwin McCravy(20056)

In this addition probleM, each letter represents the same digit in both problem. Replace each letter with a different digit, 1 through 9, so that both additon problems a true.
ABC + DEF = GHI
ADG + BEH = CFI


There are four different solutions:

Solution 1:
ABC + DEF = GHI,    ADG + BEH = CFI
146 + 583 = 729,    157 + 482 = 639

Solution 2:
ABC + DEF = GHI,    ADG + BEH = CFI
157 + 482 = 639,    146 + 583 = 729

Solution 3:
ABC + DEF = GHI,    ADG + BEH = CFI
718 + 236 = 954,    729 + 135 = 864

Solution 4:
ABC + DEF = GHI,    ADG + BEH = CFI
729 + 135 = 864,    718 + 236 = 954

Edwin